The angle of the reflected light n=2.42 whose perfectle polarized? Find the reflectance of TE and TM of that angle

1
Expert's answer
2022-02-22T04:41:25-0500

"n=2.42,"

"n_0=1,"

"n\sin\theta=n_0,"

"\theta=\arcsin\frac 1n=24.4°,"

"\theta_0=90°-\theta=65.6°."